OpenStudy (anonymous):

If f(x) = 2x2 - 30, find f(4). 35 -14 2 17

OpenStudy (uri):

Just plug in 4 instead of x and solve.

OpenStudy (anonymous):

f(4) \(f(x)=2x^2-30\\f(4)=2(4)^2-30\\f(4)=?\)

OpenStudy (uri):

2(4)^2-30 @ugh.math

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

can you help me with a couple of questions @uri ?

OpenStudy (uri):

:)

OpenStudy (uri):

Ye close this one and post in other thread.

OpenStudy (anonymous):

ok :)